# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 2.6 Averaging Samplers
The following lemma gives guarantees on sub-sampling from a weak source using an averaging sampler.
@@
$Y\circ X_{\mathsf{Samp}(Y)}\approx_{\gamma+2^{-\Omega(\tau n)}}Y\circ W,$
@@
We will need the following averaging sampler based on random walks on expanders. Let $G$ be a $D$-regular graph with vertex set $[n]$. We assume that the neighborhood of each vertex is ordered in some predetermined way. Then, the associated averaging sampler parses its input $x$ as $(i_{1},b_{1},b_{2},\dots,b_{t-1})$, where $i_{1}\in[n]$ and $b_{1},\dots,b_{t-1}\in[D]$, and outputs $\mathsf{Samp}(x)=(i_{1},\dots,i_{t})$, where $i_{j}$ is the $b_{j-1}$-th neighbor of $i_{j-1}$ when $j>1$. To ensure distinct samples, we skip repeated vertices.
The performance of $\mathsf{Samp}$ as an averaging sampler is determined by the spectral expansion of $G$. In fact, if $G$ has spectral expansion $\theta/2$ then a direct application of the expander Chernoff bound [Gil98] gives that $\mathsf{Samp}$ is a $(\gamma,\theta)$-averaging sampler with $t=O(\log(1/\gamma)/\theta^{2})$ and $r=\log n+O(t\log(1/\theta))$ [Vad04, Section 8.2]. We instantiate $G$ with the regular expander graphs from the following result of Alon [Alo21].
@@
Fix any prime $p$ such that $p\equiv 1\mod 4$. Then, there is a constant $C_{p}$ such that for every integer $n\geq C_{p}$ there exists a $(D=p+2)$-regular graph $G_{n}$ on $n$ vertices with spectral expansion $\lambda\leq\frac{(1+\sqrt{2})\sqrt{d-1}+o(1)}{d}$, where the $o(1)$ tends to 00 as $n\to\infty$. Furthermore, the family $(G_{n})_{n}$ is strongly explicit.
In particular, for any $\theta>0$ there exist constants $C_{\theta}>0$ and $D_{\theta}=O(\theta^{-2})$ and a strongly explicit family of $D_{\theta}$-regular graphs $(G_{n})_{n\geq C_{\theta}}$ with spectral expansion $\lambda\leq\theta$ for any $n\geq C_{\theta}$.
Taking the $t$-th power of a $\lambda$-spectral expander improves its expansion to $\lambda^{t}$. This readily gives us the following corollary.
@@
For every large enough $n$, and any $\lambda=\lambda(n)>0$, there exists a $D$-regular graph $G=(V=[n],E)$ with spectral expansion $\lambda$, where $D=\operatorname{poly}(1/\lambda)$, and given $x\in[n]$ and $i\in[D]$, the $y$-th neighbor of $x$ can be computed in time $\log(1/\lambda)\cdot\operatorname{polylog}(n)$.
Combining the discussion above with Lemma 2.17 (or Corollary 2.18) immediately yields the following.
@@
For every large enough integer $n$ and every $\theta,\gamma\in(0,1)$, there exists a $(\gamma,\theta)$-averaging sampler $\mathsf{Samp}\colon\{0,1\}^{r}\to[n]^{t}$ with distinct samples with $t=O(\log(1/\gamma)/\theta^{2})$ and $r=\log n+O(t\log(1/\theta))$. Furthermore, $\mathsf{Samp}$ is computable in time $O(t\cdot\operatorname{polylog}n)$.
We can extend Lemma 2.19 to output more distinct samples while not increasing $r$ via the following simple lemma.
@@
Suppose that $\mathsf{Samp}_{0}\colon\{0,1\}^{r}\to[n]^{t}$ is a $(\gamma,\theta)$-averaging sampler with distinct samples. Then, for every integer $m\geq 1$ there exists a $(\gamma,\theta)$-averaging sampler $\mathsf{Samp}\colon\{0,1\}^{r}\to[m\cdot n]^{m\cdot t}$ with distinct samples. Furthermore, if $\mathsf{Samp}_{0}$ is computable in time $T$, then $\mathsf{Samp}$ is computable in time $O(mT)$.
Lemma 2.20 follows easily by parsing $[m\cdot t]=[m]\times[t]$ and considering the sampler $\mathsf{Samp}(x)_{i,j}=(i,\mathsf{Samp}_{0}(x)_{j})$ for $i\in[m]$ and $j\in[t]$. If we can compute $\mathsf{Samp}_{0}(x)$ in time $T$, then we can compute $\mathsf{Samp}(x)$ in time $O(mT)$, as desired. The following is an easy consequence of Lemmas 2.19 and 2.20.
@@
There exists a constant $C>0$ such that the following holds. For every large enough $n$ and $\theta,\gamma\in(0,1)$, there exists a $(\gamma,\theta)$-averaging sampler $\mathsf{Samp}\colon\{0,1\}^{r}\to[n]^{t}$ with distinct samples for any $t\in[t_{0},n]$ with $t_{0}\leq C\log(1/\gamma)/\theta^{2}$ and $r=\log(n/t)+\log(1/\gamma)\cdot\operatorname{poly}(1/\theta)$. Furthermore, $\mathsf{Samp}$ is computable in time $t\cdot\operatorname{poly}(1/\theta,\log n)$.
In particular, if $\theta$ is constant then $t_{0}=O(\log(1/\gamma))$, $r=\log(n/t)+O(\log(1/\gamma))$, and $\mathsf{Samp}$ is computable in time $t\cdot\operatorname{polylog}n$.
@@ 2.7 Standard Composition Techniques for Extractors
Suppose that for $i\in\{1,2\}$ there exist strong $(k_{i},\varepsilon_{i})$-seeded extractors $\mathsf{Ext}_{i}\colon\{0,1\}^{n}\times\{0,1\}^{d_{i}}\to\{0,1\}^{m_{i}}$ running in time $T_{i}$, with $k_{2}\leq k_{1}-m_{1}-g$. Then, $\mathsf{Ext}\colon\{0,1\}^{n}\times\{0,1\}^{d_{1}+d_{2}}\to\{0,1\}^{m_{1}+m_{2}}$ given by $\mathsf{Ext}(X,(Y_{1},Y_{2}))=\mathsf{Ext}_{1}(X,Y_{1})\circ\mathsf{Ext}(X,Y_{2})$ is a strong $(k_{1},\frac{\varepsilon_{1}}{1-2^{-g}}+\varepsilon_{2})$-seeded extractor running in time $O(T_{1}+T_{2})$.
@@
Let $X=X_{1}\circ\cdots\circ X_{t}$ be an $((n_{1},\dots,n_{t}),(k_{1},\dots,k_{t}))$-block-source, and let $\mathsf{Ext}_{i}\colon\{0,1\}^{n_{i}}\times\{0,1\}^{d_{i}}\to\{0,1\}^{m_{i}}$ be average-case strong $(k_{i},\varepsilon_{i})$-seeded extractors running in time $T_{i}$ with output length $m_{i}\geq d_{i-1}-d_{i}$ for $i\geq 2$. Then, there exists a strong $(k_{1},\dots,k_{t},\varepsilon=\sum_{i\in[t]}\varepsilon_{i})$-block-source extractor $\mathsf{BExt}\colon\{0,1\}^{n_{1}}\times\cdots\times\{0,1\}^{n_{t}}\times\{0,1\}^{d_{t}}\to\{0,1\}^{m}$ with output length $m=m_{1}+\sum_{i=2}^{t}(m_{i}-(d_{i-1}-d_{i}))$ that runs in time $O(\sum_{i\in[t]}T_{i})$. If $X$ is an exact block source, then the $\mathsf{Ext}_{i}$’s do not need to be average-case.
We discuss how the fast hash-based extractor from Lemma 2.14 can be used to construct a fast extractor with seed length any constant factor smaller than its output length for high min-entropy sources. We need the following lemma, which is an easy consequence of the chain rule for min-entropy.
@@
If $X$ is an $(n,k=n-\Delta)$-source and we write $X=X_{1}\circ\cdots\circ X_{t}$ with $|X_{i}|\geq n^{\prime}$ for all $i\in[t]$, then $X_{1}\circ\cdots\circ X_{t}$ is $t\varepsilon$-close to an $(n_{1}=n^{\prime},\dots,n_{t}=n^{\prime},k^{\prime}=n^{\prime}-\Delta-\log(1/\varepsilon))$-block-source.
@@
Let $X$ be an $(n,k=(1-\frac{1}{20t})n)$-source and $\varepsilon^{\prime}=\frac{\varepsilon}{2t}$. Write $X$ as $X=X_{1}\circ\cdots\circ X_{t}$ with $|X_{i}|=\lfloor n/t\rfloor=n^{\prime}$ for all $i$. Then, Lemma 2.24 guarantees that $X_{1}\circ\cdots\circ X_{t}$ is $(t\varepsilon^{\prime})$-close to an $(n_{1}=n^{\prime},\dots,n_{t}=n^{\prime},k^{\prime}=n^{\prime}-\frac{n}{20t}-\log(1/\varepsilon^{\prime}))$-block-source $X^{\prime}$. Note that
@@
since we have assumed that $\varepsilon>2^{-\frac{n}{50t}}$. Now, let $\mathsf{Ext}^{\prime}\colon\{0,1\}^{n^{\prime}}\times\{0,1\}^{d}\to\{0,1\}^{m}$ be the strong $(k^{\prime},\varepsilon^{\prime})$-seeded extractor from Lemma 2.14 with output length $m=\mathopen{}\mathclose{{}\left\lceil\frac{k}{2t}}\right\rceil\leq k-4\log(16/\varepsilon^{\prime})$ and corresponding seed length $d\leq k/t+4\log(n/\varepsilon^{\prime})+9\leq k/t+C\log(n/\varepsilon)$ for a large enough constant $C>0$ depending only on $t$. Then, we apply block source extraction (Lemma 2.23) to $X^{\prime}$ with $\mathsf{Ext}_{1}=\cdots=\mathsf{Ext}_{t}=\mathsf{Ext}^{\prime}$ to get the desired strong $(k,2t\varepsilon^{\prime}=\varepsilon)$-extractor $\mathsf{Ext}$ with output length $t\cdot m\geq k/2$ and seed length $d$. Since $\mathsf{Ext}^{\prime}$ is computable in time $O(n\log n)$, then $\mathsf{Ext}$ is computable in time $O(tn\log n)$. ∎
In addition to Lemma 2.22, one can potentially boost the output length of a high min-entropy extractor by first treating the source as a block sources and then performing a simple block source extraction. The following corollary follows easily from Lemmas 2.23 and 2.24 (and can also be found in [Vad04, Section 6]).
@@
Let $\mathsf{Ext}_{\mathsf{in}}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n/2}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{\ell}\rightarrow\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$ and $\mathsf{Ext}_{\mathsf{out}}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n/2}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\rightarrow\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m}$ be $(k^{\prime},\varepsilon)$-extractors. Then, for any $(n,k=\delta n)$-source $X_{1}\circ X_{2}$, each $X_{i}\sim\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n/2}$, and an independent uniform $Y\sim\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{\ell}$, we have that
@@
is $4\varepsilon$-close to uniform, assuming that $k^{\prime}\geq(\delta-\frac{3}{4})n$ and $\varepsilon\geq 2^{-n/4}$. In other words, $\mathsf{Ext}$ is a $(k,4\varepsilon)$-extractor. Moreover, if $\mathsf{Ext}_{\mathsf{in}}$ is strong then $\mathsf{Ext}$ is also strong, and if $\mathsf{Ext}_{\mathsf{in}}$ and $\mathsf{Ext}_{\mathsf{out}}$ run in time $T_{1}$ and $T_{2}$, respectively, then $\mathsf{Ext}$ runs in time $O(T_{1}+T_{2})$.

### v2

@@ 2.6 Averaging Samplers
It is worth noting the equivalence between averaging samplers and extractors [ZUC97], where we think of the samples obtained by enumerating over the seeds of the extractor. Specifically, a $(k,\varepsilon)$-seeded extractor over inputs of length $r$ is a $(2^{k-r+1},\varepsilon)$-averaging sampler, and a $(\gamma,\theta)$-averaging sampler with input length $r$ is a $(k,2\theta)$-seeded extractor for $k=n-\log(1/\gamma)+\log(1/\theta)$.
The following lemma gives guarantees on sub-sampling coordinates from a weak source using an averaging sampler.
@@
$\mathopen{}\mathclose{{\left(Y,X_{\mathsf{Samp}(Y)}}}\right)\approx_{\gamma+2^{-\Omega(\tau n)}}(Y,W),$
@@
We will need the following well-known averaging sampler based on random walks on expanders (see, e.g., [GIL98, ZUC07]). Let $G$ be a $D$-regular graph with vertex set $[n]$. We assume that the neighborhood of each vertex is ordered in some predetermined way. Then, the associated averaging sampler parses its input $x$ as $(i_{1},b_{1},b_{2},\dots,b_{t-1})$, where $i_{1}\in[n]$ and $b_{1},\dots,b_{t-1}\in[D]$, and outputs $\mathsf{Samp}(x)=(i_{1},\dots,i_{t})$, where $i_{j}$ is the $b_{j-1}$-th neighbor of $i_{j-1}$ when $j>1$.
The performance of $\mathsf{Samp}$ as an averaging sampler is determined by the spectral expansion of GG. [^12] We say that an undirected graph $G$ over $n$ vertices has spectral expansion $\lambda$ if $\lambda\leq\max_{i\geq 2}|\lambda_{i}|$, where $\lambda_{n}\leq\ldots\leq\lambda_{2}\leq\lambda_{1}=1$ are the eigenvalues of $G$’s random walk matrix. In fact, if $G$ has spectral expansion $\theta/2$ then a direct application of the expander Chernoff bound [GIL98] gives that $\mathsf{Samp}$ is a $(\gamma,\theta)$-averaging sampler with $t=O(\log(1/\gamma)/\theta^{2})$ and $r=\log n+O(t\log(1/\theta))$ [VAD04, Section 8.2].
To ensure distinct samples while maintaining accuracy, we follow [VAD04] and modify the standard random walk sampler as follows. As a “base” sampler, we use the above walk, but only take the first $(1-\theta/2)t$ distinct vertices (if there are such). Letting $r_{0}=\log n+O(t\log(1/\theta))$ be the corresponding randomness complexity, we then take a random walk of length $\ell=O(\log(1/\gamma))$ over an expander $G_{0}$ with $2^{r_{0}}$ vertices and constant spectral expansion. Each vertex of $G_{0}$ corresponds to a random walk on $G$, and out of those $\ell$ chosen walks, we take the first one that indeed has $(1-\theta/2)t$ distinct vertices (and output some arbitrary value if the process failed). In [VAD04, Lemma 8.2], it is shown that setting the spectral expansion of $G$ to be $\Theta(\theta)$ (and suitably choosing the constants inside the $O()$ notation), one still gets a $(\gamma,\theta)$-averaging sampler.
We instantiate $G$ and $G_{0}$ with the regular expander graphs from the following result of Alon [ALO21].
@@
Fix any prime $p$ such that $p\equiv 1\mod 4$. Then, there is a constant $C_{p}$ such that for every integer $n\geq C_{p}$ there exists a $(D=p+2)$-regular graph $G_{n}$ on $n$ vertices with spectral expansion $\lambda\leq\frac{(1+\sqrt{2})\sqrt{D-1}+o(1)}{D}$, where the $o(1)$ tends to 0 as $n\to\infty$. Furthermore, the family $(G_{n})_{n}$ is strongly explicit.
In particular, for any $\theta>0$ there exist constants $C_{\theta}>0$ and $D_{\theta}=O(\theta^{-2})$ and a strongly explicit family of $D_{\theta}$-regular graphs $(G_{n})_{n}$ with spectral expansion $\lambda\leq\theta$ for any $n\geq C_{\theta}$.
Given a graph $G$, its $t$-th power $G^{t}$ is a graph over the same number of vertices, and each edge corresponds to a $t$-step walk over $G$. It is well-known that taking the $t$-th power of a $\lambda$-spectral expander improves its expansion to $\lambda^{t}$. This readily gives us the following corollary.
@@
For every large enough $n$, and any $\lambda=\lambda(n)>0$, there exists a $D$-regular graph $G=(V=[n],E)$ with spectral expansion $\lambda$, where $D=\operatorname{poly}(1/\lambda)$, and given $x\in[n]$ and $i\in[D]$, the $i$-th neighbor of $x$ can be computed in time $\log(1/\lambda)\cdot\operatorname{polylog}(n)=\operatorname{polylog}(n)$.
Combining the discussion above with lemma˜2.16 (or corollary˜2.17) immediately yields the following, observing that the runtime is dominated by $\ell\cdot t\log(1/\lambda)\operatorname{polylog}(n)$.
@@
For every large enough integer $n$ and every $\theta,\gamma\in(0,1)$, there exists a $(\gamma,\theta)$-averaging sampler $\mathsf{Samp}\colon\{0,1\}^{r}\to[n]^{t}$ with distinct samples with $t=O(\log(1/\gamma)/\theta^{2})$ and $r=\log n+O(t\log(1/\theta))$. Furthermore, $\mathsf{Samp}$ is computable in time $t\log(1/\gamma)\cdot\operatorname{polylog}n$.
We can extend lemma˜2.18 to output more distinct samples while not increasing $r$ via the following simple lemma.
@@
Suppose that $\mathsf{Samp}_{0}\colon\{0,1\}^{r}\to[n]^{t}$ is a $(\gamma,\theta)$-averaging sampler with distinct samples. Then, for every integer $m\geq 1$ there exists a $(\gamma,\theta)$-averaging sampler $\mathsf{Samp}\colon\{0,1\}^{r}\to[m\cdot n]^{m\cdot t}$ with distinct samples.
lemma˜2.19 follows easily by parsing $m\cdot t]=[m]\times[t]$ and considering the sampler $\mathsf{Samp}(x)_{i,j}=(i,\mathsf{Samp}_{0}(x)_{j})$ for $i\in[m]$ and $j\in[t]$. If we can compute $\mathsf{Samp}_{0}(x)$ in time $T$, then we can compute $\mathsf{Samp}(x)$ in time $T+O(mt\log(mn))$. The following is an easy consequence of lemmas˜2.18 and 2.19.
@@
There exists a constant $C>0$ such that the following holds. For every large enough $n$ and $\theta,\gamma\in(0,1)$, there exists a $(\gamma,\theta)$-averaging sampler $\mathsf{Samp}\colon\{0,1\}^{r}\to[n]^{t}$ with distinct samples for any $t\in[t_{0},n]$ with $t_{0}\leq C\log(1/\gamma)/\theta^{2}$ and $r=\log(n/t)+\log(1/\gamma)\cdot\operatorname{poly}(1/\theta)$. Furthermore, $\mathsf{Samp}$ is computable in time $t_{0}\log(1/\gamma)\cdot\operatorname{polylog}n+O(t\log n)$.
In particular, if $\theta$ is constant then $t_{0}=O(\log(1/\gamma))$, $r=\log(n/t)+O(\log(1/\gamma))$, and $\mathsf{Samp}$ is computable in time $\log^{2}(1/\gamma)\cdot\operatorname{polylog}n+O(t\log n)$.
@@ 2.7 Standard Composition Techniques for Extractors
Suppose that for $i\in\{1,2\}$ there exist strong $(k_{i},\varepsilon_{i})$-seeded extractors $\mathsf{Ext}_{i}\colon\{0,1\}^{n}\times\{0,1\}^{d_{i}}\to\{0,1\}^{m_{i}}$ running in time $T_{i}$, with $k_{2}\leq k_{1}-m_{1}-g$. Then, $\mathsf{Ext}\colon\{0,1\}^{n}\times\{0,1\}^{d_{1}+d_{2}}\to\{0,1\}^{m_{1}+m_{2}}$ given by $\mathsf{Ext}(X,(Y_{1},Y_{2}))=(\mathsf{Ext}_{1}(X,Y_{1}),\mathsf{Ext}_{2}(X,Y_{2}))$ is a strong $(k_{1},\frac{\varepsilon_{1}}{1-2^{-g}}+\varepsilon_{2})$-seeded extractor running in time $O(T_{1}+T_{2})$.
@@
Let $X=(X_{1},\dots,X_{t})$ be an $((n_{1},\dots,n_{t}),(k_{1},\dots,k_{t}))$-block-source, and let $\mathsf{Ext}_{i}\colon\{0,1\}^{n_{i}}\times\{0,1\}^{d_{i}}\to\{0,1\}^{m_{i}}$ be average-case $(k_{i},\varepsilon_{i})$-seeded extractors running in time $T_{i}$ with output length $m_{i}\geq d_{i-1}$ for $i\geq 2$, that output their seed. Then, there exists a strong $(k_{1},\dots,k_{t},\varepsilon=\sum_{i\in[t]}\varepsilon_{i})$-block-source extractor $\mathsf{BExt}\colon\{0,1\}^{n_{1}}\times\cdots\times\{0,1\}^{n_{t}}\times\{0,1\}^{d_{t}}\to\{0,1\}^{m}$ with output length $m=m_{1}-d_{t}$ that runs in time $O(\sum_{i\in[t]}T_{i})$. If $X$ is an exact block source, then the $\mathsf{Ext}_{i}$’s do not need to be average-case.
We discuss how the fast hash-based extractor from lemma˜2.13 can be used to construct a fast extractor with seed length any constant factor smaller than its output length for high min-entropy sources. We need the following lemma, which is an easy consequence of the chain rule for min-entropy.
@@
Let $X$ be an $(n,k=n-\Delta)$-source and let $X_{1},\dots,X_{t}$ correspond to disjoint subsets of coordinates of $X$, with each $X_{i}$ of length $n_{i}\geq n^{\prime}$. Then, $(X_{1},\dots,X_{t})$ is $t\varepsilon$-close to an exact $((n_{1},\dots,n_{t}),(k^{\prime},\dots,k^{\prime}))$-block-source for $k^{\prime}=n^{\prime}-\Delta-\log(1/\varepsilon)$.131313[GUV09, Corollary 4.16] is originally stated only for the special case where $X_{1},\dots,X_{t}$ partition the coordinates of $X$. This extends easily to the statement presented here by noting that the puncturing of $X$ to its $\widetilde{n}=\sum_{i=1}^{t}n_{i}$ bits corresponding to $X_{1},\dots,X_{t}$ is an $(\widetilde{n},\widetilde{n}-\Delta)$-source, and applying the original statement in [GUV09].
@@
Let $X$ be an $(n,k=(1-\frac{1}{20t})n)$-source and $\varepsilon^{\prime}=\frac{\varepsilon}{2t}$. Let $X_{1},\dots,X_{t}$ correspond to disjoint subsets of coordinates of $X$, with $|X_{i}|=\lfloor n/t\rfloor=n^{\prime}$ for all $i$. Then, lemma˜2.23 guarantees that $(X_{1},\ldots,X_{t})$ is $(t\varepsilon^{\prime})$-close to an exact $((n_{1}=n^{\prime},\dots,n_{t}=n^{\prime}),k^{\prime}=n^{\prime}-\frac{n}{20t}-\log(1/\varepsilon^{\prime}))$-block-source $X^{\prime}$. Note that
@@
since we have assumed that $\varepsilon>2^{-\frac{n}{50t}}$. Now, let $\mathsf{Ext}^{\prime}\colon\{0,1\}^{n^{\prime}}\times\{0,1\}^{d}\to\{0,1\}^{m}$ be the strong $(k^{\prime},\varepsilon^{\prime})$-seeded extractor from lemma˜2.13 with output length $m=\mathopen{}\mathclose{{\left\lceil\frac{k}{2t}}}\right\rceil\leq k-4\log(16/\varepsilon^{\prime})$ and corresponding seed length $d\leq k/t+4\log(n^{\prime}/\varepsilon^{\prime})+9\leq k/t+C\log(n/\varepsilon)$ for a large enough constant $C>0$ depending only on $t$. Then, we apply block source extraction (lemma˜2.22) to $X^{\prime}$ with $\mathsf{Ext}_{1}=\cdots=\mathsf{Ext}_{t}=\mathsf{Ext}^{\prime}$ to get the desired strong $(k,2t\varepsilon^{\prime}=\varepsilon)$-extractor $\mathsf{Ext}$ with output length $t\cdot m\geq k/2$ and seed length $d$. Since $\mathsf{Ext}^{\prime}$ is computable in time $O(n\log n)$, then $\mathsf{Ext}$ is computable in time $O(tn\log n)$. ∎
In addition to lemma˜2.21, one can potentially boost the output length of a high min-entropy extractor by first treating the source as a block source and then performing a simple block source extraction. The next corollary appears in [VAD12, Lemma 6.27].
@@
Let $\mathsf{Ext}_{\mathsf{in}}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n/2}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\ell}\rightarrow\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$ and $\mathsf{Ext}_{\mathsf{out}}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n/2}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\rightarrow\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m}$ be $(k^{\prime},\varepsilon)$-extractors. Then, for any $(n,k=\delta n)$-source $(X_{1},X_{2})$, each $X_{i}\sim\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n/2}$, and an independent uniform $Y\sim\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\ell}$, we have that
@@
is $4\varepsilon$-close to uniform, assuming that $k^{\prime}\leq(\delta-\frac{3}{4})n$ and $\varepsilon\geq 2^{-n/4}$. In other words, $\mathsf{Ext}$ is a $(k,4\varepsilon)$-extractor. Moreover, if $\mathsf{Ext}_{\mathsf{in}}$ is strong then $\mathsf{Ext}$ is also strong, and if $\mathsf{Ext}_{\mathsf{in}}$ and $\mathsf{Ext}_{\mathsf{out}}$ run in time $T_{1}$ and $T_{2}$, respectively, then $\mathsf{Ext}$ runs in time $O(T_{1}+T_{2})$.